Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH UK

Tel: 0703 593081

email: dbc@uk.ac.soton.ecs

October 26, 1995

Prof. Geoffrey C. Fox, Director, NPAC Syracuse University New York 13244-4100 USA

Dear Professor Fox,

With reference to your email of 25th October, 1995, thank you for offering me a position as Research Scientist at NPAC, starting January 1996.

I am very happy to accept this offer, according to the conditions stated in your letter.

I look forward to seeing you in the new year,

yours sincerely,